### v1

@@ 1 Introduction
Meanwhile, strong magnetic fields are expected to be created in relativistic heavy-ion collisions at Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) and also in astrophysical objects such as magnetars Harding:2006qn ; Hattori:2016emy , motivating extensive studies of hadrons and QCD matter under intense magnetic fields (see Refs. Miransky:2015ava ; Hattori:2023egw ; Adhikari:2024bfa for review). Notable examples include the (inverse) magnetic catalysis and its impact on chiral symmetry Gusynin:1994xp ; DElia:2011koc ; Bali:2012zg ; Fukushima:2012kc ; Bruckmann:2013oba ; DElia:2018xwo , as well as modifications of hadron spectra and their internal structures under strong fields Alford:2013jva ; Taya:2014nha ; Hattori:2015aki ; Hattori:2019ijy .
@@
$\mathcal{L}=$ $\frac{f_{\pi}^{2}}{4}\eta^{\mu\nu}\mathop{\mathrm{tr}}\left(D_{\mu}UD_{\nu}U^{{\dagger}}\right)+\frac{f_{\pi}^{2}\mathcal{B}}{2}\mathop{\mathrm{tr}}\left(MU^{\dagger}+UM^{\dagger}\right)$
$+{\bar{\mathcal{N}}}\left[\mathrm{i}\gamma^{\mu}D_{\mu}-m_{N}+\frac{1}{2}g_{A}\gamma^{\mu}\gamma_{5}\widetilde{\alpha}_{\mu\perp}\right]\mathcal{N},$ (1)
@@
with the pion mass $m_{\pi}:=2m\mathcal{B}$, the spin Pauli matrices $\sigma^{i}~(i=1,2,3)$, and the covariant derivative for nucleons,
@@
Note that the covariant derivative contains the gauge field $A_{\mu}$, through which the $NN\pi$ vertex acquires the dependence on the external gauge field. We use an isospin basis defined by
@@
As a final remark, we clarify the power counting scheme employed in this paper. We count the momentum $p_{\mu}$ of pions and nonrelativistic nucleons and the background gauge field $eA_{\mu}$ as $\mathcal{O}(\epsilon)$ quantities, while the quark-mass matrix is counted as $M=\mathcal{O}(\epsilon^{2})$. This counting allows us to take magnetic-field strengths as large as the pion-mass scale, $|e\bm{B}|\sim m_{\pi}^{2}=\mathcal{O}(\epsilon^{2})$. As a rough estimate of the ultraviolet cutoff $\Lambda_{\mathrm{UV}}$, we take $\Lambda_{\mathrm{UV}}\simeq m_{\rho}=770\,$MeV. Accordingly, we numerically consider magnetic-field strengths up to $|e\bm{B}|\lesssim m_{\rho}^{2}\sim 30m_{\pi}^{2}$. [^2] This estimate is admittedly crude. For instance, it neglects the mass shifts of other hadrons due to the applied magnetic field Adhikari:2024bfa . A more realistic estimate would take into account the mass shift of, for example, the $\rho$ meson, which we do not consider here.
@@ 3.1 Deriving two-nucleon interaction Hamiltonian
where the unsummed indice $a=0,\pm$ denotes the electric charge of the pion field, and $\tau_{-a}$ is the corresponding isospin Pauli matrix, as defined in Eq. (6). In the following analysis, we focus on the static case and neglect the temporal derivatives of the pion field. Introducing a coordinate-space Green’s function for pions in the presence of an external gauge field, $G^{a}({\bm{x},\bm{x}^{\prime}}|A)$, such that
@@ 3.2 Definition of gauge-invariant potential
The Hamiltonian (16) encodes all the information about the interaction between two nucleons induced by the one-pion exchange. To define the OPEP, there are two possible approaches: (i) fixing a specific gauge [e.g., the Fock–Schwinger (FS) gauge] and (ii) constructing a gauge-invariant OPEP by dressing the nucleon field with a spatial Wilson line. The second approach was adopted in Ref. Deshmukh:2018owh . Although physical observables derived from the potential, such as binding energies, are gauge invariant, it is conceptually advantageous to formulate the potential itself in a manifestly gauge-invariant way.
Let us describe the approach (ii) and its connection to the approach (i) in more detail. We choose a reference position $\bm{x}_{0}$ and introduce a “dressed” nucleon field $\widetilde{N}(\bm{x})$ as
@@ 3.3 Weak and strong magnetic-field limits
where $\bm{r}=:(\bm{r}_{\perp},z)$ and $r=|\bm{r}|$. We note that the first line of Eq. (29) exactly conincides with $V_{0}$, after evaluating the derivatives in Eq. (26), with the tensor operator,
@@
When the applied magnetic field is sufficiently strong compared with the pion mass in vacuum ($|eB|\gg m_{\pi}^{2}$), the charged pions undergo the Landau quantization, and the lowest energy mode, or the lowest Landau level (LLL), dominates. In this regime, the charged-pion Green’s function can be approximated well by the contribution from the LLL [i.e., the $n=0$ mode in Eq. (47)] as
$\begin{split}G_{B}({\bm{x}-\bm{x}^{\prime}})\simeq\frac{|eB|}{4\pi}\mathrm{e}^{-\frac{|eB|}{4}|{\bm{x}_{\perp}-\bm{x}_{\perp}^{\prime}}|^{2}}\frac{\mathrm{e}^{-\sqrt{m_{\pi}^{2}+|eB|}|z-z^{\prime}|}}{\sqrt{m_{\pi}^{2}+|eB|}}.\end{split}$ (31)
With this, the magnetic-field dependence in the strong-field limit can be computed as
$\begin{split}{\lim_{eB\gg m_{\pi}^{2}}}V_{pn}(\bm{r})&={\lim_{eB\gg m_{\pi}^{2}}}\left[V_{np}(\bm{r})\right]^{\dagger}\\ &=-\frac{g_{A}^{2}|eB|}{16\pi f_{\pi}^{2}}\Bigg[\frac{|eB|}{2\sqrt{m_{\pi}^{2}+|eB|}}\bm{\sigma}^{(1)}_{\perp}\cdot\bm{\sigma}^{(2)}_{\perp}-\frac{|eB|^{2}}{4\sqrt{m_{\pi}^{2}+|eB|}}(\bm{\sigma}^{(1)}_{\perp}\cdot\bm{r}_{\perp})(\bm{\sigma}^{(2)}_{\perp}\cdot\bm{r}_{\perp})\\ &\hskip 60.0pt-\frac{|eB|}{2}\left(\sigma^{3}_{(1)}(\bm{\sigma}^{(2)}_{\perp}\cdot\bm{r}_{\perp})+(\bm{\sigma}^{(1)}_{\perp}\cdot\bm{r}_{\perp})\sigma^{3}_{(2)}\right)\frac{z}{|z|}-\sqrt{m_{\pi}^{2}+|eB|}\sigma^{3}_{(1)}\sigma^{3}_{(2)}\\ &\hskip 70.0pt+\frac{|eB|^{2}}{4\sqrt{m_{\pi}^{2}+|eB|}}\left(\bm{\sigma}_{(1)}\times\bm{r}_{\perp}\right)_{z}\left(\bm{\sigma}_{(2)}\times\bm{r}_{\perp}\right)_{z}\Bigg]\mathrm{e}^{-\frac{|eB|}{4}\bm{r}_{\perp}^{2}-\sqrt{m_{\pi}^{2}+|eB|}|z|}.\end{split}$ (32)
@@ 4.1 OPEP in the isospin singlet channel
On the other hand, in Fig. 4, we compare the full result with the lowest-Landau-level (LLL) approximation given in Eq. (32) at $|eB|=25m_{\pi}^{2}$. As one can clearly see, the LLL contribution dominates in the long-distance region (e.g., $z\gtrsim 3$ fm and $r_{\perp}\gtrsim 2$ fm in the plot) because the LLL corresponds to the lightest excitation mode. Higher Landau levels are heavier and therefore contribute only at shorter distances, which explains the observed deviations between the LLL approximation and the full result at small $z$ and $r_{\perp}$.
@@ 5 Energy shift of deuteron
For the unperturbed deuteron state, we use the known one based on the AV18 potential Wiringa:1994wb . [^4] We note that even if the unperturbed eigenstates in a magnetic field are defined as the dressed ones [i.e., with the Wilson lines (17)], the matrix elements $A_{MM^{\prime}}$ can still be evaluated using the undressed states, because the Wilson lines from the bra and ket cancel each other, i.e., $\langle{\tilde{d}_{M}}|\big(V_{\text{OPEP}}^{B\neq 0}-V_{\text{OPEP}}^{B=0}\big)|{\tilde{d}_{M^{\prime}}}\rangle=\langle{d_{M}}|\big(V_{\text{OPEP}}^{B\neq 0}-V_{\text{OPEP}}^{B=0}\big)|{d_{M^{\prime}}}\rangle.$
@@ 6 Summary and discussion
First, additional magnetic-field corrections to the OPEP or nuclear forces should be taken into account, such as the Zeeman energy shift and contributions from heavy mesons—particularly the charged $\rho$-meson, whose mass may decrease in magnetic fields. These are also important to discuss the deuteron, or the nucleon bound-state problem in general. As an example, consider the Zeeman energy shift. A rough estimate gives that the resulting energy shift is the order of $E_{\mathrm{Zeeman}}\sim eB/2m_{N}\sim 10\,\mathrm{MeV}$ at $|eB|\sim m_{\pi}^{2}$, which is much larger than the deuteron’s binding energy $E^{B=0}\simeq-2.2\,$MeV. Therefore, to obtain more realistic understanding of nuclear forces and their physical impacts in strong magnetic fields of order $|eB|\sim m_{\pi}^{2}$, these additional mangetic-field effects must be incorporated in future studies.
@@ Acknowledgment
The authors thank Koichi Hattori and Yoshimasa Hidaka for stimulating discussion. D.M. is supported by JST SPRING, Grant No. JPMJSP2121 and JST, the establishment of University fellowships towards the creation of science technology innovation, Grant No. JPMJFS2114. M.H. is supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grants No. JP23K25870, No. 25K01002, and No. JP25K07316. H.T. is supported by JSPS KAKENHI Grant No. 24K17058 and the RIKEN TRIP initiative (RIKEN Quantum). T.H. was partly supported by Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2318. This work was partially supported by the RIKEN iTHEMS.

### v2

@@ 1 Introduction
Meanwhile, strong magnetic fields are expected to be created in relativistic heavy-ion collisions at Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) and also in astrophysical objects such as magnetars Harding:2006qn ; Hattori:2016emy , motivating extensive studies of hadrons and QCD matter under intense magnetic fields (see Refs. Miransky:2015ava ; Hattori:2023egw ; Adhikari:2024bfa for review). Notable examples include the (inverse) magnetic catalysis and its impact on chiral symmetry Gusynin:1994xp ; DElia:2011koc ; Bali:2012zg ; Fukushima:2012kc ; Bruckmann:2013oba ; DElia:2018xwo , as well as modifications of hadron spectra and their internal structures under strong fields Tiburzi:2008ma ; Alford:2013jva ; Taya:2014nha ; Hattori:2015aki ; Bali:2017ian ; Hattori:2019ijy .
@@
$\mathcal{L}=$ $\frac{f_{\pi}^{2}}{4}\eta^{\mu\nu}\mathop{\mathrm{tr}}\left(D_{\mu}UD_{\nu}U^{{\dagger}}\right)+\frac{f_{\pi}^{2}\mathcal{B}}{2}\mathop{\mathrm{tr}}\left(M^{\dagger}U+U^{\dagger}M\right)$
$+{\bar{\mathcal{N}}}\left[\mathrm{i}\gamma^{\mu}D_{\mu}-m_{N}+g_{A}\gamma^{\mu}\gamma_{5}\widetilde{\alpha}_{\mu\perp}\right]\mathcal{N},$ (1)
@@
with the pion mass squared $m_{\pi}^{2}:=2m\mathcal{B}$, the spin Pauli matrices $\sigma^{i}~(i=1,2,3)$, and the covariant derivative for nucleons,
@@
Note that the covariant derivative contains the gauge field $A_{\mu}$, through which the $NN\pi$ vertex acquires the dependence on the external gauge field. We use an isospin basis defined by [^1] With our normalization in Eq. (6), the product of isospin operators is decomposed as $\bm{\tau}^{(1)}\!\cdot\!\bm{\tau}^{(2)}=\tau^{(1)}_{+}\tau^{(2)}_{-}+\tau^{(1)}_{-}\tau^{(2)}_{+}+\tau^{(1)}_{0}\tau^{(2)}_{0}$. It differs from another common convention $\tau^{\prime}_{\pm}:=(\tau_{1}\pm\mathrm{i}\tau_{2})/2$, for which $\bm{\tau}^{(1)}\!\cdot\!\bm{\tau}^{(2)}=2\tau^{(1)\prime}_{+}\tau^{(2)\prime}_{-}+2\tau^{(1)\prime}_{-}\tau^{(2)\prime}_{+}+\tau^{(1)}_{0}\tau^{(2)}_{0}$.
@@
As a final remark, we clarify the power counting scheme employed in this paper. We count the momentum $p_{\mu}$ of pions and nonrelativistic nucleons and the background gauge field $eA_{\mu}$ as $p_{\mu}=\mathcal{O}(\epsilon)$ and $eA_{\mu}=\mathcal{O}(\epsilon)$, respectively, while the quark-mass matrix is counted as $M=\mathcal{O}(\epsilon^{2})$. This counting allows us to take magnetic-field strengths as large as the pion-mass scale, $|e\bm{B}|\sim m_{\pi}^{2}=\mathcal{O}(\epsilon^{2})$. In this scheme, higher-order terms of $\mathcal{O}(\epsilon^{4})$ or above lie beyond the accuracy of our leading-order analysis and are neglected. As a rough estimate of the ultraviolet cutoff $\Lambda_{\mathrm{UV}}$, we take $\Lambda_{\mathrm{UV}}\simeq m_{\rho}=770\,$MeV. Accordingly, we numerically consider magnetic-field strengths up to $|e\bm{B}|\lesssim m_{\rho}^{2}\sim 30m_{\pi}^{2}$. [^3] This estimate is admittedly crude. For instance, it neglects the mass shifts of other hadrons due to the applied magnetic field Adhikari:2024bfa . A more realistic estimate would take into account the mass shift of, for example, the $\rho$ meson, which we do not consider here. For magnetic fields approaching the upper end of this range, higher-order contributions beyond $\mathcal{O}(\epsilon^{2})$ may become quantitatively important, and the present leading-order analysis should be regarded as subject to truncation uncertainties of the chiral expansion.
@@ 3.1 Deriving two-nucleon interaction Hamiltonian
where the unsummed indices $a=0,\pm$ denote the electric charge of the pion field, and $\tau_{-a}$ is the corresponding isospin Pauli matrix, as defined in Eq. (6). In the following analysis, we focus on the static case and neglect the temporal derivatives of the pion field. Introducing a coordinate-space Green’s function for pions in the presence of an external gauge field, $G^{a}({\bm{x},\bm{x}^{\prime}}|A)$, such that
@@ 3.2 Definition of gauge-invariant potential
The Hamiltonian (16) encodes all the information about the interaction between two nucleons induced by the one-pion exchange. Physical observables calculated from this Hamiltonian, such as scattering cross sections and binding energies, are gauge independent. On the other hand, the OPEP itself is not an observable; it is an intermediate, gauge-dependent quantity, whose definition requires a specific prescription.
In principle, any gauge choice can be used to define such a potential as long as observables remain unaffected. In this work, we adopt a manifestly gauge-invariant definition by introducing the Wilson line Deshmukh:2018owh . The Wilson line cancels the gauge-dependent Schwinger phase associated with the charged-pion propagator, ensuring that the same potential is obtained in any gauge. In particular, in the Fock–Schwinger (FS) gauge the Wilson line becomes unity, so the potential can be evaluated without explicit reference to the Schwinger phase. This prescription allows us to visualize the OPEP as a unique function of the relative coordinate, which should facilitate comparisons with other approaches.
Let us now detail the formulation and demonstrate that this gauge-invariant construction reduces to the expression in the FS gauge. We choose a reference position $\bm{x}_{0}$ and introduce a “dressed” nucleon field $\widetilde{N}(\bm{x})$ as
@@ 3.3 Weak and strong magnetic-field limits
where $\bm{r}=:(\bm{r}_{\perp},z)$ and $r=|\bm{r}|$. We note that the first line of Eq. (29) exactly coincides with $V_{0}$, after evaluating the derivatives in Eq. (26), with the tensor operator,
@@
When the applied magnetic field is sufficiently strong, the dynamics of charged pions is dominated by the lowest Landau level (LLL), and a similar LLL dominance is expected to appear in the OPEP. To clarify the conditions for this dominance, we recall that the charged-pion dynamics becomes anisotropic in a magnetic field.
The longitudinal dynamics is governed by the effective mass $\sqrt{m_{\pi}^{2}+(2n+1)|eB|}$ of the Landau levels ($n=0,1,2,\cdots$), whereas the transverse dynamics is controlled by the magnetic length $1/\sqrt{|eB|}$, which sets the size of the Landau orbit [see Eqs. (43)–(48) in Appendix A]. As a result, the LLL ($n=0$) becomes dominant along the $z$-direction when the longitudinal separation satisfies $|z|\gg 1/\sqrt{m_{\pi}^{2}+|eB|}$. In contrast, the LLL dominance in the transverse direction arises when $|\bm{r}_{\perp}|\gg 1/\sqrt{|eB|}$. In the strong-field regime $|eB|\gg m_{\pi}^{2}$, the longitudinal condition also reduces to $|z|\gg 1/\sqrt{|eB|}$.
In the LLL-dominant regime, the charged-pion Green’s function can be approximated well by the contribution from the LLL [i.e., the $n=0$ mode in Eq. (47)] as
$\begin{split}G_{B}({\bm{x}-\bm{x}^{\prime}})\approx G_{B}^{\mathrm{LLL}}({\bm{x}-\bm{x}^{\prime}}):=\frac{|eB|}{4\pi}\mathrm{e}^{-\frac{|eB|}{4}|{\bm{x}_{\perp}-\bm{x}_{\perp}^{\prime}}|^{2}}\frac{\mathrm{e}^{-\sqrt{m_{\pi}^{2}+|eB|}|z-z^{\prime}|}}{\sqrt{m_{\pi}^{2}+|eB|}}.\end{split}$ (31)
Substituting the LLL Green’s function $G_{B}^{\mathrm{LLL}}(\bm{x}-\bm{x}^{\prime})$ into Eq. (26), the OPEP induced by charged-pion exchange can be written in the LLL-dominant regime as
$\begin{split}V_{pn}^{\text{LLL}}(\bm{r})&=\left[V_{np}^{\text{LLL}}(\bm{r})\right]^{\dagger}\\ &=-\frac{g_{A}^{2}|eB|}{16\pi f_{\pi}^{2}}\Bigg[\frac{|eB|}{2\sqrt{m_{\pi}^{2}+|eB|}}\bm{\sigma}^{(1)}_{\perp}\cdot\bm{\sigma}^{(2)}_{\perp}-\frac{|eB|^{2}}{4\sqrt{m_{\pi}^{2}+|eB|}}(\bm{\sigma}^{(1)}_{\perp}\cdot\bm{r}_{\perp})(\bm{\sigma}^{(2)}_{\perp}\cdot\bm{r}_{\perp})\\ &\hskip 60.0pt-\frac{|eB|}{2}\left(\sigma^{3}_{(1)}(\bm{\sigma}^{(2)}_{\perp}\cdot\bm{r}_{\perp})+(\bm{\sigma}^{(1)}_{\perp}\cdot\bm{r}_{\perp})\sigma^{3}_{(2)}\right)\frac{z}{|z|}-\sqrt{m_{\pi}^{2}+|eB|}\sigma^{3}_{(1)}\sigma^{3}_{(2)}\\ &\hskip 70.0pt+\frac{|eB|^{2}}{4\sqrt{m_{\pi}^{2}+|eB|}}\left(\bm{\sigma}_{(1)}\times\bm{r}_{\perp}\right)_{z}\left(\bm{\sigma}_{(2)}\times\bm{r}_{\perp}\right)_{z}\Bigg]\mathrm{e}^{-\frac{|eB|}{4}\bm{r}_{\perp}^{2}-\sqrt{m_{\pi}^{2}+|eB|}|z|}.\end{split}$ (32)
@@ 4.1 OPEP in the isospin singlet channel
On the other hand, in Fig. 4, we compare the full result with the lowest-Landau-level (LLL) approximation given in Eq. (32) at $|eB|=25m_{\pi}^{2}$. As one can clearly see, the LLL contribution dominates in the long-distance region (e.g., $r_{\perp}\gtrsim 2$ fm in the plot). This behavior is consistent with the condition for LLL dominance, $r_{\perp}\gg 1/\sqrt{|eB|}$. At $|eB|=25m_{\pi}^{2}$, the magnetic length is $1/\sqrt{|eB|}=1/(5m_{\pi})\approx 0.3$ fm. Therefore, at short transverse distances $r_{\perp}\lesssim 1/\sqrt{|eB|}$, higher-Landau-level contributions become non-negligible, and the LLL approximation deviates from the full result.
@@ 5 Energy shift of deuteron
For the unperturbed deuteron state (i.e., the wave function in the absence of the magnetic field), we use the wave function obtained from the AV18 potential, which includes both the $^{3}S_{1}$ and $^{3}D_{1}$ components. [^5] We note that even if the unperturbed eigenstates in a magnetic field are defined as the dressed ones [i.e., with the Wilson lines (17)], the matrix elements $A_{MM^{\prime}}$ can still be evaluated using the undressed states, because the Wilson lines from the bra and ket cancel each other, i.e., $\langle{\tilde{d}_{M}}|\big(V_{\text{OPEP}}^{B\neq 0}-V_{\text{OPEP}}^{B=0}\big)|{\tilde{d}_{M^{\prime}}}\rangle=\langle{d_{M}}|\big(V_{\text{OPEP}}^{B\neq 0}-V_{\text{OPEP}}^{B=0}\big)|{d_{M^{\prime}}}\rangle.$
@@ 6 Summary and discussion
First, additional magnetic-field effects that are not included in the present analysis should be taken into account. For instance, the Zeeman energy shift is expected to be large when discussing the deuteron binding energy or nucleon bound states in general. A rough estimate gives that the resulting energy shift is the order of $E_{\mathrm{Zeeman}}\sim eB/2m_{N}\sim 10\,\mathrm{MeV}$ at $|eB|\sim m_{\pi}^{2}$, which is much larger than the deuteron’s binding energy $E^{B=0}\simeq-2.2\,$MeV. Therefore, to obtain more realistic understanding of nuclear forces and their physical impacts in strong magnetic fields of order $|eB|\sim m_{\pi}^{2}$, these additional magnetic-field effects must be incorporated in future studies.
In addition, strong magnetic fields modify the masses of mesons mediating the nuclear force. For example, the charged $\rho$-meson mass is expected to decrease with increasing magnetic field, and various theoretical studies also indicate that the neutral- and charged-pion masses change in strong magnetic fields (see, e.g., Refs. Taya:2014nha ; Bali:2017ian ). Such modifications in the meson spectrum would feed back into the meson-exchange contributions to the nuclear force, including the OPEP. Inclusion of these effects is an important direction toward a more quantitative description of nuclear forces in strong magnetic fields.
@@ Acknowledgment
The authors thank Koichi Hattori and Yoshimasa Hidaka for stimulating discussion. D.M. is supported by JST SPRING, Grant No. JPMJSP2121 and JST, the establishment of University fellowships towards the creation of science technology innovation, Grant No. JPMJFS2114. M.H. is supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grants No. 23K25870, No. 25K01002, and No. 25K07316. H.T. is supported by JSPS KAKENHI Grant No. 24K17058 and the RIKEN TRIP initiative (RIKEN Quantum). T.H. was partly supported by Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2318. This work was partially supported by the RIKEN iTHEMS and Niigata University Quantum Research Center (NU-Q).